1,3,7,15,31, ,127 what comes after 31?
Answer by rothauserc(4718)

1, 3, 7, 15, 31, ,127
This is an arithmetic sequence whose nth term is defined as
a(subscript n) = a(subscript n-1) + 2^(n-1)
a6 = 31 + 2^5
a6 = 31 + 32 = 63
therefore 63 comes after 31